a=(2i+3j) , b=(4i-2j) and c=(1i+4j) evaluate the volume of the shape
1
Expert's answer
2020-10-13T14:44:22-0400

Solution

Given vectors are

a=(2i+3j)

b=(4i-2j)

c=(1i+4j)

Now volume of shape is given by

"V=a.(b\u00d7c)" ................ eq. 1

Now

Cross product of vector b and c

(b×c)=18 k

And

a.(b×c)=(2i+3j). (18k) =0

Now putting these value in equation 1

Volume of given shape

V=0